Asset GBP

Budgeted profits for the year 10,575,000

Asset base at the start of the year 58,925,000

Budgeted asset base at the end of the year 69,500,000

Average asset base 64,212,500

ROA (average of opening and closing assets) 16.5%

Calculate the budgeted ROA for sales outputs of 4,200,000 and 4,500,000 units.
1
Expert's answer
2020-05-20T10:12:11-0400

Return on assets is a profitability ratio that provides how much profit a company is able to generate from its assets. In other words, ROA measures how efficient a company's management is in generating earnings from their economic resources or assets on their balance sheet. ROA is shown as a percentage, and the higher the number, the more efficient a company's management is at managing its balance sheet to generate profits.


"ROA=Net Income\/Average Total Assets*100"


ROA1=4200000/64212500*100=6,54%

RQA2=4500000/64212500*100=7,00%



https://www.investopedia.com/ask/answers/031215/what-formula-calculating-return-assets-roa.asp